OpenStudy (anonymous):

Evaluate 2 - 3m - m^2 when m = -3

OpenStudy (anonymous):

please work it out step by step because I dont know what to do first!

OpenStudy (anonymous):

-m^2 - 3m + 2 -(-3)^2 -3(-3) + 2 3^2 = 3*3 = 9 -3(-3) = -3*-3 = 9 9 + 9 + 2 = 20 I believe.

OpenStudy (anonymous):

My teacher says the answer is 2

OpenStudy (anonymous):

Ahaha... time for a reevaluation on my part... O_O sorry :b

OpenStudy (anonymous):

Its okay!!